# Exact Mobility Edges in a Disorder-Free Dimerized Stark Lattice with Effective Unbounded Hopping

### v1

@@ II Model and Exact Solution
We consider a single-particle Hamiltonian on a quasi-1D Creutz ladder [as shown in Fig. 1(a)] with $L=2N+1$ unit cells, indexed by $n\in\{-N,\dots,N\}$. Each unit cell contains two sublattices, AA and $B$. The real-space Schrödinger equations governing the system are given by
@@
To obtain the bulk spectrum in the thermodynamic limit, we transition to reciprocal space using the Fourier transform. We start from the original Hamiltonian Eq. 1 without transformation. In this representation, the position operator transforms as $n\to-i\partial_{k}$ [Hartmann_2004]. The Schrödinger equation becomes a system of coupled first-order differential equations
@@
$\arg[\phi_{B}(k)]$ $=\frac{t_{1}^{2}+t_{2}^{2}-E^{2}}{Ft_{2}\sqrt{1-\alpha^{2}}}\ln\left(\frac{\alpha\tan(k/2)-1-\sqrt{1-\alpha^{2}}}{\alpha\tan(k/2)-1+\sqrt{1-\alpha^{2}}}\right)$ (6)
$-\frac{2t_{1}}{F}\ln(\alpha-\sin k),$
@@
$\Theta(k)\sim-\frac{\text{Res}(k_{c})}{F}\ln|k-k_{c}|,\quad\text{Res}(k_{c})=\frac{(t_{1}+t_{2}\cos k_{c})^{2}}{t_{2}\cos k_{c}}.$ (11)
@@ Appendix B Restrictions on the Spectrum of Jacobi Matrices
Next, we examine broader conditions for the existence or absence of spectral components in unbounded Jacobi matrices in mathematical physics. We will demonstrate that our effective Hamiltonian $H^{\prime\prime}$ does not satisfy the sufficient conditions for an empty AC spectrum or empty pure point ($P$) spectrum found in the relevant literature [Cojuhari_2008, Swiderski_2016].
@@
Regarding the $P$ spectrum, Ref. [Swiderski_2016] [Theorem 1.1] provides conditions forbidding the $P$ spectrum of a Jacobi matrix. A necessary condition for this theorem to hold is
@@
For our model, taking the subsequence $a_{2n}=t_{2}$, we have $|b_{2n}|/a_{2n}\propto n/t_{2}\to\infty$. Thus, the theorem does not apply, and the existence of a $P$ spectrum (our localized Stark ladder) is allowed.
@@ Appendix C Derivation of the Spectrum
$I(k)$ $=\int^{k}\frac{E^{2}-t_{1}^{2}-t_{2}^{2}-2t_{1}t_{2}\cos q}{E-t_{2}\sin q}dq$ (27)
$=(E^{2}-t_{1}^{2}-t_{2}^{2})\int^{k}\frac{dq}{E-t_{2}\sin q}+\int^{k}\frac{2t_{1}t_{2}\cos q}{E-t_{2}\sin q}dq.$
@@
$\phi_{B}(k)=\left(\frac{\alpha\tan\frac{k}{2}-1-\sqrt{1-\alpha^{2}}}{\alpha\tan\frac{k}{2}-1+\sqrt{1-\alpha^{2}}}\right)^{-\frac{i(E^{2}-t_{1}^{2}-t_{2}^{2})}{Ft_{2}\sqrt{1-\alpha^{2}}}}\cdot(\alpha-\sin k)^{-2it_{1}/F}.$ (29)

### v2

@@ II Model and Exact Solution
We consider a single-particle Hamiltonian on a quasi-1D Creutz ladder [as shown in Fig. 1(a)] with $L=2N+1$ unit cells, indexed by $n\in\{-N,\dots,N\}$. Each unit cell contains two sublattices, $A$ and $B$. The real-space Schrödinger equations governing the system are given by
@@
To obtain the bulk spectrum in the thermodynamic limit, we transition to reciprocal space using the Fourier transform. We start from the original Hamiltonian Eq. 1 without transformation. In this representation, the position operator transforms as $n\to i\partial_{k}$ [Hartmann_2004]. The Schrödinger equation becomes a system of coupled first-order differential equations
@@
$\arg[\phi_{B}(k)]$ $=\frac{E^{2}-t_{1}^{2}-t_{2}^{2}}{Ft_{2}\sqrt{1-\alpha^{2}}}\ln\left(\frac{\alpha\tan(k/2)-1-\sqrt{1-\alpha^{2}}}{\alpha\tan(k/2)-1+\sqrt{1-\alpha^{2}}}\right)$ (6)
$+\frac{2t_{1}}{F}\ln(\alpha-\sin k),$
@@
$\Theta(k)\sim-\frac{\text{Res}(k_{c})}{F}\ln|k-k_{c}|,\quad\text{Res}(k_{c})=-\frac{(t_{1}+t_{2}\cos k_{c})^{2}}{t_{2}\cos k_{c}}.$ (11)
@@ Appendix B Restrictions on the Spectrum of Jacobi Matrices
Next, we examine broader conditions for the existence or absence of spectral components in unbounded Jacobi matrices in mathematical physics. We will demonstrate that our effective Hamiltonian $H^{\prime\prime}$ does not satisfy the sufficient conditions for an empty AC spectrum or empty pure point (PP) spectrum found in the relevant literature [Cojuhari_2008, Swiderski_2016].
@@
Regarding the PP spectrum, Ref. [Swiderski_2016] [Theorem 1.1] provides conditions forbidding the PP spectrum of a Jacobi matrix. A necessary condition for this theorem to hold is
@@
For our model, taking the subsequence $a_{2n}=t_{2}$, we have $|b_{2n}|/a_{2n}\propto n/t_{2}\to\infty$. Thus, the theorem does not apply, and the existence of a PP spectrum (our localized Stark ladder) is allowed.
@@ Appendix C Derivation of the Spectrum
$I(k)$ $=\int^{k}\frac{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos q-E^{2}}{E-t_{2}\sin q}dq$ (27)
$=(t_{1}^{2}+t_{2}^{2}-E^{2})\int^{k}\frac{dq}{E-t_{2}\sin q}+\int^{k}\frac{2t_{1}t_{2}\cos q}{E-t_{2}\sin q}dq.$
@@
$\phi_{B}(k)=\left(\frac{\alpha\tan\frac{k}{2}-1-\sqrt{1-\alpha^{2}}}{\alpha\tan\frac{k}{2}-1+\sqrt{1-\alpha^{2}}}\right)^{\frac{i(E^{2}-t_{1}^{2}-t_{2}^{2})}{Ft_{2}\sqrt{1-\alpha^{2}}}}\cdot(\alpha-\sin k)^{2it_{1}/F}.$ (29)